OpenStudy (sgayton27):

Find the domain of the function sqrt(4+x)/3-x

OpenStudy (sgayton27):

\[\frac{ \sqrt{4+x} }{3-x }\]

OpenStudy (sgayton27):

i understand that x cannot equal 3. but i cannot get the correct answer. I thought the answer was \[(-\infty,3)U(3,\infty)\] but its not

OpenStudy (lauluisa):